# Theoretical study on the crystal structure of a bilayer nickel-oxychloride Sr3Ni2O5Cl2 and analysis on the occurrence of possible unconventional superconductivity

### v2

@@ I Introduction
Discovery of superconductivity with a maximum $T_{c}$ of above 80 K in a bilayer Ruddlesden-Popper nickelate La3Ni2O7 (Fig. 1(a)) under pressure has sparked enormous interest in the field of condensed matter physics [1]. Although under pressure, $T_{c}$ exceeding the liquid nitrogen boiling temperature, and also the origin of superconductivity likely being unconventional [2, 3, 4], have led to a huge amount of experimental and theoretical studies after the discovery [1, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 3, 4, 69, 2, 70, 71, 72, 73, 74, 75, 76, 77]. In fact, the possibility of superconductivity in this material was discussed theoretically even before its experimental discovery by some of the previous authors [78], but one viewpoint that was missing at that time was the precise crystal structure, namely, although a tetragonal symmetry was simply assumed in Ref. 78, the actual structural symmetry is orthorhombic at ambient pressure. There are now accumulating studies suggesting that tetragonal symmetry of the crystal structure may be important for the occurrence of superconductivity in La3Ni2O7 [32, 42, 41], which is accomplished by applying pressure. One speculation regarding the reason for this can be that orthorhombicity favors density waves, which suppresses superconductivity. Interestingly, also for the trilayer compound La4Ni3O10, the superconducting phase appears by applying a sufficient pressure and the tetragonal structure is realized [31, 38, 39, 40, 48]. Then, a natural motivation arises for searching materials that have tetragonal symmetry at ambient pressure, and also have electronic properties similar to that of La3Ni2O7 at high pressures.
@@ II Methods
For first-principles calculation, we used the PBEsol exchange-correlation functional [83] and the projector augmented wave method as implemented in Vienna ab initio Simulation Package (VASP) [84, 85, 86, 87, 88]. Core-electron states in PAW potentials were [Kr]$4d^{10}$, [Ar]$3d^{10}$, [Ar], [Ne], and [He] for La, Sr, Ni, Cl, and O, respectively. We used a plane-wave cutoff energy of 600 eV for Kohn-Sham orbitals without including the spin-orbit coupling for simplicity. A $12\times 12\times 12$ ${\bm{k}}$-mesh was used for self-consistent-field calculations.
@@ III.1 Electronic and phonon band structures for Sr3Ni2O5Cl2
As mentioned above, the average valence number of the nickel atom is different: Ni2.5+ ($d^{7.5}$) for La3Ni2O7 and Ni3+ ($d^{7}$) for Sr3Ni2O5Cl2. Thus, the number of $d$ electrons for Sr3Ni2O5Cl2 is less than that for La3Ni2O7. Nevertheless, the Ni-$d_{3z^{2}-r^{2}}$ orbital is rather closer to the half-filling in Sr3Ni2O5Cl2 as shown in Figs. 2(b)(h). This is because of the characteristic crystal field for the nickel atom in Sr3Ni2O5Cl2, where a half of apical O2- ions is replaced with a Cl- ion, which pushes down the orbital energy of the Ni-$d_{3z^{2}-r^{2}}$ orbital. As a result, the number of $d_{x^{2}-y^{2}}$ electrons is decreased in Sr3Ni2O5Cl2. On the other hand, Ni-$d_{xy/yz/zx}$ orbitals, which we hereafter call them the $t_{2g}$ orbitals while it is not a rigorous term for this symmetry, raise their energy levels relative to the Ni-$d_{3z^{2}-r^{2}}$ orbital. This is because the present crystal field strongly pushes down the Ni-$d_{3z^{2}-r^{2}}$ orbital level as we have stated above. Consequently, the strong DOS peak of the $t_{2g}$ orbitals lies just below the Fermi energy. In addition, the crystal field is strongly asymmetric with respect to the $xy$ plane (i.e., $+z$ and $-z$ are strongly inequivalent), which allows a sizable hybridization among $e_{g}$ ($d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$) and $t_{2g}$ orbitals. In the next section, we will discuss how these characteristics can affect superconductivity in this system.
@@ III.2 Model construction
Table 1: The orbital level offset $\Delta E=E_{x^{2}-y^{2}}-E_{3z^{2}-r^{2}}$ between the $d_{x^{2}-y^{2}}$ and the $d_{3z^{2}-r^{2}}$ orbitals, the vertical interlayer hopping $t_{\perp}$ between the $d_{3z^{2}-r^{2}}$ orbitals, and the nearest-neighbor intralayer hoppings $t_{3z^{2}-r^{2}}$, $t_{x^{2}-y^{2}}$, and $t_{x^{2}-y^{2}{\text{-}}3z^{2}-r^{2}}$ are displayed. The first column indicates the chemical formulae of the calculated materials, where the number in the round brackets indicate the number of orbitals.
@@
In Fig. 3(b), we plot the number of electrons for each orbital per Ni against the hypothetically varied total band filling, assuming a rigid band. In this paper, the total band filling $n$ is defined as the number of electrons per spin in the primitive unit cell, e.g., $0\leq n\leq 10$ for ten-orbital models. The doping ratio $\Delta n$ is defined as the deviation of the total band filling from its value at stoichiometry. In Sr3Ni2O5Cl2, the number of electrons occupying the $d_{3z^{2}-r^{2}}$ band per Ni, called $n[d_{3z^{2}-r^{2}}]$ hereafter, exceeds unity (half-filled) already at stoichiometry, consistent with the observation given in Section III.1. Also, the $d_{yz/zx}$ orbitals are not fully filled due to the hybridization with the $e_{g}$ orbitals. Contrastingly, in La3Ni2O7, the $d_{yz/zx}$ orbitals are fully filled in the whole doping range suggesting their irrelevance for superconductivity and justifying the previously proposed four-orbital model [8].
@@ III.4 Formation energy (enthalpy)
Therefore, we can expect that the material we proposed in this study, Sr3Ni2O5Cl2, can be synthesized with a help of the high-pressure synthesis. Our calculation also suggests that, under sufficiently high pressure, Sr2NiO2Cl2 can appear as impurity phases because the formation enthalpy for chemical reaction (c) is not effectively lowered by pressure as shown in Fig. 5. This result suggests that Sr2NiO2Cl2 is possibly included as an impurity phase in the high-pressure synthesis of Sr3Ni2O5Cl2 under some conditions.

### v3

# Theoretical study of the crystal structure of the bilayer nickel oxychloride Sr3Ni2O5Cl2 and analysis of possible unconventional superconductivity

@@ I Introduction
Discovery of superconductivity with a maximum $T_{c}$ of above 80 K in a bilayer Ruddlesden-Popper nickelate La3Ni2O7 [Fig. 1(a)] under pressure has sparked enormous interest in the field of condensed matter physics [1]. Although under pressure, $T_{c}$ exceeding the liquid nitrogen boiling temperature, and also the origin of superconductivity likely being unconventional [2, 3, 4], have led to a huge amount of experimental and theoretical studies after the discovery [1, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 3, 4, 69, 2, 70, 71, 72, 73, 74, 75, 76, 77]. In fact, the possibility of superconductivity in this material was discussed theoretically even before its experimental discovery by some of the previous authors [78], but one viewpoint that was missing at that time was the precise crystal structure, namely, although a tetragonal symmetry was simply assumed in Ref. 78, the actual structural symmetry is orthorhombic at ambient pressure. There are now accumulating studies suggesting that tetragonal symmetry of the crystal structure may be important for the occurrence of superconductivity in La3Ni2O7 [32, 42, 41], which is accomplished by applying pressure. One speculation regarding the reason for this can be that orthorhombicity favors density waves, which suppresses superconductivity. Interestingly, also for the trilayer compound La4Ni3O10, the superconducting phase appears by applying a sufficient pressure and the tetragonal structure is realized [31, 38, 39, 40, 48]. Then, a natural motivation arises for searching materials that have tetragonal symmetry at ambient pressure, and also have electronic properties similar to that of La3Ni2O7 at high pressures.
@@ II Methods
For first-principles calculation, we used the PBEsol exchange-correlation functional [83] and the projector augmented wave (PAW) method as implemented in Vienna ab initio Simulation Package (VASP) [84, 85, 86, 87, 88]. Core-electron states in PAW potentials were [Kr]$4d^{10}$, [Ar]$3d^{10}$, [Ar], [Ne], and [He] for La, Sr, Ni, Cl, and O, respectively. We used a plane-wave cutoff energy of 600 eV for Kohn-Sham orbitals without including the spin-orbit coupling for simplicity. A $12\times 12\times 12$ ${\bm{k}}$-mesh was used for self-consistent-field calculations.
@@ III.1 Electronic and phonon band structures for Sr3Ni2O5Cl2
As mentioned above, the average valence number of the nickel atom is different: Ni2.5+ ($d^{7.5}$) for La3Ni2O7 and Ni3+ ($d^{7}$) for Sr3Ni2O5Cl2. Thus, the number of $d$ electrons for Sr3Ni2O5Cl2 is less than that for La3Ni2O7. Nevertheless, the Ni-$d_{3z^{2}-r^{2}}$ orbital is rather closer to the half-filling in Sr3Ni2O5Cl2 as shown in Figs. 2(b)(h). This is because of the characteristic crystal field for the nickel atom in Sr3Ni2O5Cl2, where a half of apical O2- ions is replaced with a Cl- ion, which pushes down the orbital energy of the Ni-$d_{3z^{2}-r^{2}}$ orbital. As a result, the number of $d_{x^{2}-y^{2}}$ electrons is decreased in Sr3Ni2O5Cl2. On the other hand, Ni-$d_{xy/yz/zx}$ orbitals, which we hereafter call the $t_{2g}$ orbitals while it is not a rigorous term for this symmetry, raise their energy levels relative to the Ni-$d_{3z^{2}-r^{2}}$ orbital. This is because the present crystal field strongly pushes down the Ni-$d_{3z^{2}-r^{2}}$ orbital level as we have stated above. Consequently, the strong DOS peak of the $t_{2g}$ orbitals lies just below the Fermi energy. In addition, the crystal field is strongly asymmetric with respect to the $xy$ plane (i.e., $+z$ and $-z$ are strongly inequivalent), which allows a sizable hybridization among $e_{g}$ ($d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$) and $t_{2g}$ orbitals. In the next section, we will discuss how these characteristics can affect superconductivity in this system.
@@ III.2 Model construction
Table 1: The orbital level offset $\Delta E=E_{x^{2}-y^{2}}-E_{3z^{2}-r^{2}}$ between the $d_{x^{2}-y^{2}}$ and the $d_{3z^{2}-r^{2}}$ orbitals, the vertical interlayer hopping $t_{\perp}$ between the $d_{3z^{2}-r^{2}}$ orbitals, and the nearest-neighbor intralayer hoppings $t_{3z^{2}-r^{2}}$, $t_{x^{2}-y^{2}}$, and $t_{x^{2}-y^{2}{\text{-}}3z^{2}-r^{2}}$ are displayed. The first column indicates the chemical formulas of the calculated materials, where the number in the round brackets indicate the number of orbitals.
@@
In Fig. 3(b), we plot the number of electrons for each orbital per Ni against the hypothetically varied total band filling, assuming a rigid band. In this paper, the total band filling $n$ is defined as the number of electrons per spin in the primitive unit cell, e.g., $0\leq n\leq 10$ for ten-orbital models. The doping ratio $\Delta n$ is defined as the deviation of the total band filling from its value at stoichiometry. In Sr3Ni2O5Cl2, the number of electrons occupying the $d_{3z^{2}-r^{2}}$ band per Ni, called $n[d_{3z^{2}-r^{2}}]$ hereafter, exceeds unity (half-filled) already at stoichiometry, consistent with the observation given in Sec. III.1. Also, the $d_{yz/zx}$ orbitals are not fully filled due to the hybridization with the $e_{g}$ orbitals. Contrastingly, in La3Ni2O7, the $d_{yz/zx}$ orbitals are fully filled in the whole doping range suggesting their irrelevance for superconductivity and justifying the previously proposed four-orbital model [8].
@@ III.4 Formation energy (enthalpy)
Therefore, we can expect that the material we proposed in this study, Sr3Ni2O5Cl2, can be synthesized with the help of the high-pressure synthesis. Our calculation also suggests that, under sufficiently high pressure, Sr2NiO2Cl2 can appear as impurity phases because the formation enthalpy for chemical reaction (c) is not effectively lowered by pressure as shown in Fig. 5. This result suggests that Sr2NiO2Cl2 is possibly included as an impurity phase in the high-pressure synthesis of Sr3Ni2O5Cl2 under some conditions.